# PLGSLAM: Progressive Neural Scene Represenation with Local to Global Bundle Adjustment

### v1

@@ Abstract
Neural implicit scene representations have recently shown encouraging results in dense visual SLAM. However, existing methods produce low-quality scene reconstruction and low-accuracy localization performance when scaling up to large indoor scenes and long sequences. These limitations are mainly due to their single, global radiance field with finite capacity, which does not adapt to large scenarios. Their end-to-end pose networks are also not robust enough with the growth of cumulative errors in large scenes. To this end, we present PLGSLAM, a neural visual SLAM system which performs high-fidelity surface reconstruction and robust camera tracking in real time. To handle large-scale indoor scenes, PLGSLAM proposes a progressive scene representation method which dynamically allocates new local scene representation trained with frames within a local sliding window. This allows us to scale up to larger indoor scenes and improves robustness (even under pose drifts). In local scene representation, PLGSLAM utilizes tri-planes for local high-frequency features. We also incorporate multi-layer perceptron (MLP) networks for the low-frequency feature, smoothness, and scene completion in unobserved areas. Moreover, we propose local-to-global bundle adjustment method with a global keyframe database to address the increased pose drifts on long sequences. Experimental results demonstrate that PLGSLAM achieves state-of-the-art scene reconstruction results and tracking performance across various datasets and scenarios (both in small and large-scale indoor environments). The code will be open-sourced upon paper acceptance.
@@ 1 Introduction
Visual Simultaneous Localization and Mapping (SLAM) has been a fundamental computer vision problem with wide applications such as autonomous driving, robotics, and virtual/augmented reality. Many traditional methods have been introduced in the past years, such as DTAM [14], ORB-SLAM [13, 12], and VINS [16]. They can estimate the camera pose and construct sparse point cloud maps in real-time with accurate localization performance. However, the sparse point cloud maps cannot meet the further perception needs of the robot. Recent attention has turned to learning-based methods for dense scene reconstruction. Kinectfusion [7], BAD-SLAM[17] reconstruct meaningful global 3D maps and show reasonable but limited reconstruction accuracy with deep learning networks.
Nowadays, with the proposal of Neural Radiance Fields (NeRF), many works focus on combining implicit scene representation with SLAM systems. iMAP [19] is the first work to use a single multi-layer perceptron (MLP) to represent the entire scene. NICE-SLAM [29] improves the scene representation method with feature grids. ESLAM [8] and Co-SLAM [23] further improve the scene representation methods. ESLAM uses tri-planes for better real-time performance and reconstruction accuracy. Co-SLAM uses joint coordinate and sparse parametric scene for accurate scene representation. They can achieve promising reconstruction quality in a small indoor room.
@@
In local scene representation, we combine the tri-planes with the multi-layer perceptron (MLP) to better represent the scene. We use tri-planes to encode the local high-frequency feature of the scene and use MLP to represent global low-frequency features with the coherence priors inherent. We bring together the benefits of both methods for accuracy, smoothness, and hole-filling in areas without observation.
@@
  • A progressive scene representation method is proposed which dynamically initiate local scene representation trained with frames within a local window. This enables scalability to large indoor scenes and long videos and significantly enhances the robustness.
  • In local scene representation, We design a joint tri-planes and multi-layer perceptron scene representation method for accurate and smooth surface reconstruction. It can not only enhance the ability of scene representation, but also substantially reduce the memory growth from cubic to square.
  • We integrate the traditional SLAM system with an end-to-end pose estimation network. A local-to-global bundle adjustment algorithm is proposed, which can mitigate cumulative error in large-scale indoor scenes. Our system manages the local and global keyframe database with the system operation, enabling bundle adjustment across all past observations, from local to global.
@@ 2 Related Work
Dense Visual SLAM. SLAM has been an active field for the past two decades. Traditional visual SLAM algorithms [12, 16, 9] estimate accurate camera poses and use sparse point clouds as the map representation. They use manipulated key points for tracking, mapping, relocalization, and loop closing. Dense visual SLAM approaches focus on reconstructing a dense map of a scene. DTAM [14] is one of the pioneer works that use the dense map and view-centric scene representation. KinectFusion [7] performs camera tracking via projective iterative-closest-point (ICP) and explicitly represents the surface of the environment via TSDF-Fusion. Some works [17, 5, 21] propose bundle adjustment(BA) method to optimize keyframe poses and construct the dense 3D structure jointly. In contrast to previous SLAM approaches, we adopt implicit scene representation of the geometry and directly optimize them during mapping.
Implicit Scene Representation.  With the proposal of Neural radiance fields (NeRF) [11], many researchers explore taking the advantages of the implicit method into 3D reconstruction. NeRF is a ground-breaking method for novel view synthesis using differentiable rendering. However, the representation of volume densities can not commit the geometric consistency, leading to poor surface prediction for reconstruction tasks. In order to deal with it, UNISURF [15] and NeuS [24] are proposed, combining world-centric 3D geometry representation with neural radiance fields. They replace the volume density with Signed Distance Field (SDF) values. Other methods [1, 2, 3, 20] use various scene geometry representation methods, such as truncated signed distance function, voxel grid, or occupancy grid with latent codes. For large-scale representation, Mega-NeRF and LocalRF [22, 10] use multiple local scene representations for the entire scene.
NeRF-based SLAM.  Some works focus on pose estimation of NeRF, iNeRF [28], NeRF- - [25] estimates the camera pose with inverse NeRF optimization when the neural implicit network is fully trained. However, those methods can not optimize poses and neural implicit network simultaneously. iMAP [19] and NICE-SLAM [29] are successively proposed to combine neural implicit mapping with SLAM. iMAP uses a single multi-layer perceptron (MLP) to represent the scene, and NICE-SLAM uses a learnable hierarchical feature grid. Co-SLAM [23] and ESLAM [8] further improves the scene representation method. These are the works most relevant to our approach. However, all of them have difficult in large-scale indoor environments and long sequences. With the proposed progressive scene representation method, we can scale up to larger indoor scenarios. The fusion of tri-planes and MLP leads to high-fidelity and smooth surface reconstruction in local scene. A local-to-global bundle adjustment method is also proposed. This method can effectively eliminate growing cumulative errors in existing methods in large indoor scenes.
@@ 3 Method
The pipeline of our system is shown in Fig. 1. We use a set of sequential RGB-D frames $\{I_{i},D_{i}\}_{i=1}^{M}$ with known camera intrinsic $K\in R_{3\times 3}$ as our input. Our model predicts camera poses $\{R_{i}|t_{i}\}^{M}_{i=1}$, color $\mathbf{c}$, and an implicit truncated signed distance (TSDF) representation $\phi_{\boldsymbol{g}}$ that can be used in marching cubes algorithm to extract 3D meshes. For the implicit mapping thread, a progressive scene representation method (Sec. 3.1) is designed to represent large-scale indoor environments. Then, in the local radiance fields, we improve the scene representation methods and combine the tri-planes with multi-layer perceptron (MLP) by our designed architecture. Sec. 3.2 walks through the rendering process, which converts raw representations into pixel depths, colors, and sdf values. For the camera tracking thread, a local-to-global bundle adjustment method (Sec. 3.3) is designed for robust and accurate pose estimation. Several carefully designed loss functions are proposed to jointly optimize the scene implicit representation and camera pose estimation. The network is incrementally updated with the system operation.
@@ 3.1 Progressive Scene Representation
where $SR_{\theta_{n}}^{n}$ denotes the local scene representation. Each local scene representation is centered at the position of the last estimated camera pose. We supervise each scene representation with a local subset of the frames. Each subset contains some overlap frames, which is important for achieving consistent reconstructions in the local scene representation. Whenever the estimated camera pose leaves the bound of the current scene representation, we stop optimizing previous ones (freeze the network parameters). At this point, we can reduce memory requirements by removing unnecessary supervisory frames. We also stop updating the mapping parameters in the tracking thread to reduce errors. If the estimated camera pose is outside the current bounds, but within a previous local scene representation, we activate the previous one and proceed with the optimization process. We further increase the global consistency by inverse distance weight (IDW) fusion for all overlapping scene representations at any supervising frame.
Local Scene Representation. Voxel grid-based architectures [6, 29, 23] are the mainstream in NeRF-based SLAM system. However, they struggle with cubical memory growing and real-time performance. Inspired by [8], we design a tri-plane architecture joint with multi-layer perceptron (MLP). We store and optimize high-frequency features(e.g. texture) on perpendicular axis-aligned planes. The MLPs are used to encode and store low-frequency features for the coherence and smoothness priors. This joint scene representation architecture achieve high-fidelity and smoothness scene reconstruction with the ability of hole filling.
Specifically, the tri-feature planes are at two scales, i.e., coarse and fine. Coarse-level representation allows efficient reconstruction of free space with fewer sample points and optimization iterations. Then the tri-planes feature $\boldsymbol{T}(x)$ can be formulated as:
@@
where $\gamma(\mathbf{x})$ represents coordinate position encoding. We use one-blob position encoding [23] instead of embedding spatial coordinates into multiple frequency bands. Finally, the color decoder predicts the RGB value:
@@
$\phi_{\boldsymbol{a}}(x)$ represents the color of the sample points. Combining the MLP with the tri-planes scene representation, our architecture achieve accurate and smooth surface reconstruction, efficient memory use, and hole filling performance.
@@ 3.2 Differentiable Rendering
where $\beta\in\mathbb{R}$ is a learnable parameter that controls the sharpness of the surface boundary. Then we define the termination probability, depth, and color as:
@@
Depth-guided Sampling. The depth images provide valuable geometry information which can guide neural point sampling along a ray within the bounds of depth uncertainty. we get $N$ points from stratified sampling between the near and far planes. Then, $N_{surface}$ points are drawn from near-surface points within the range $[d-\Delta d,d+\Delta d]$, where $\Delta d$ is a small offset. If the pixel does not have a valid corresponding depth value, we use an estimated depth instead.
@@ 3.3 Local-to-global Bundle Adjustment
Currently, existing nerf-slam methods exhibit poor accuracy and robustness in large-scale indoor scene localization. Their tracking networks are performed via minimizing rgb loss functions with respect to learnable parameters $\theta$ to estimate the relative pose matrix $\{R_{i}|t_{i}\}\in\mathbb{SE}(3)$. With the growing cumulative error $\varepsilon$ of pose estimation, those methods result in failure in large-scale indoor scenes and long videos. To this end, we design a local-to-global bundle adjustment method to solve this problem, which performs well with our progressive scene representation. We design our method by drawing inspiration from traditional keyframe-based SLAM systems for improving the robustness and accuracy of pose estimation. We propose neural warpping error and reprojection error for local-to-global bundle adjustment. The neural warpping loss is formulated as:
@@
Whenever a keyframe arrives, we perform local bundle adjustment in our tracking and mapping thread. A keyframe is selected for every $K$ frames. When the camera moves to the bound of the current scene representation, we also initialize it as the keyframe. In local bundle adjustment, we only select keyframes from the local keyframe database that visually overlap with the current frame when optimizing the scene geometry to ensure the geometry outside the current view remains static and fast convergence. Meanwhile, we also maintain a global keyframe list with the operation of our system. After accumulating a specific number of local keyframes or the camera moves to the local bound, a global bundle adjustment is performed. In global BA, we randomly select keyframes and rays from the global keyframe database, which leverages all historical observations of the scene. This approach effectively integrates local and global information which greatly improves the robustness and accuracy of camera pose optimization in large-scale indoor scenes.
@@
Table 1: Quantitative results of our proposed PLGSLAM with existing NeRF-SLAM system on the Replica dataset [18]. We evaluate reconstruction and localization performance in small room scenes. The results are the average on the scenes of the Replica dataset. Our method outperforms the existing method in surface reconstruction and pose estimation.
@@ 3.4 Objective Functions
Ours 6.033 11.086 3.261 86.557 5.574 6.228
Table 3: Quantitative results of our proposed PLGSLAM with existing NeRF-SLAM system on the Apartment dataset [29]. We evaluate reconstruction and localization performance in large-scale multi-room scenes. The results are the average of three runs. Our method outperforms the existing method in surface reconstruction and pose estimation.
@@ 4.1 Datasets and Metrics
Figure 5: Completion ratio vs. model size and average time for PLGSLAM with other methods. Each model corresponds to a different hash-table size. iMAP and NICE-SLAM shown for reference.
@@
Implementation We run PLGSLAM on a desktop PC with NVIDIA RTX 3090ti GPU. For experimental settings, the truncation distance $tr$ is set to 6 cm. We employ feature planes with a resolution of 24 cm for coarse tri-planes. We use 6cm resolution for fine tri-planes. All feature planes have 32 channels, resulting in a 64-channel concatenated feature input for the decoders. The decoders are two-layer MLPs with 32 channels in the hidden layer. For Replica [18], we sample $N$ = 32 points for stratified sampling and $N_{surface}$ = 8 points for importance sampling on each ray. And for ScanNet [4], we set $N$ = 48 and $N_{surface}$ = 8. We set $\lambda_{fs}=5$ , $\lambda_{sdf_{m}}=200$ ,$\lambda_{sdf_{t}}=10$ , $\lambda_{smooth}=0.01$, $\lambda_{d}=0.1$ , and $\lambda_{c}=5$ And during tracking, we set $\lambda_{fs}=10$ , $\lambda_{sdf_{m}}=200$, $\lambda_{sdf_{t}}=50$ , $\lambda_{re}=10$ , $\lambda_{nw}=10$ , $\lambda_{d}=1$, and $\lambda_{c}=5$ For further details of our implementation, refer to the supplementary.

### v2

@@ Abstract
Neural implicit scene representations have recently shown encouraging results in dense visual SLAM. However, existing methods produce low-quality scene reconstruction and low-accuracy localization performance when scaling up to large indoor scenes and long sequences. These limitations are mainly due to their single, global radiance field with finite capacity, which does not adapt to large scenarios. Their end-to-end pose networks are also not robust enough with the growth of cumulative errors in large scenes. To this end, we introduce PLGSLAM, a neural visual SLAM system capable of high-fidelity surface reconstruction and robust camera tracking in real-time. To handle large-scale indoor scenes, PLGSLAM proposes a progressive scene representation method which dynamically allocates new local scene representation trained with frames within a local sliding window. This allows us to scale up to larger indoor scenes and improves robustness (even under pose drifts). In local scene representation, PLGSLAM utilizes tri-planes for local high-frequency features with multi-layer perceptron (MLP) networks for the low-frequency feature, achieving smoothness and scene completion in unobserved areas. Moreover, we propose local-to-global bundle adjustment method with a global keyframe database to address the increased pose drifts on long sequences. Experimental results demonstrate that PLGSLAM achieves state-of-the-art scene reconstruction results and tracking performance across various datasets and scenarios (both in small and large-scale indoor environments).
@@ 1 Introduction
Visual Simultaneous Localization and Mapping (SLAM) has been a fundamental computer vision problem with wide applications such as autonomous driving, remote sensing [6], and virtual/augmented reality. Many traditional methods have been introduced in the past years, such as ORB-SLAM [20, 19], VINS [24], and so on. They can estimate the camera pose and construct sparse point cloud maps in real-time with accurate localization performance. However, the sparse point cloud maps cannot meet the further perception needs of the robot. Recent attention has turned to learning-based methods for dense scene reconstruction. Kinectfusion [11], BAD-SLAM[25] reconstruct meaningful global 3D maps and show reasonable but limited reconstruction accuracy with deep learning networks.
Nowadays, with the proposal of Neural Radiance Fields (NeRF), there are many following works on different areas [8]. iMAP [27] is the first work to use a single multi-layer perceptron (MLP) to represent the entire scene in SLAM system. NICE-SLAM [41] improves the scene representation method with feature grids. ESLAM [12] and Co-SLAM [31] further improve the scene representation methods. ESLAM uses tri-planes for better real-time performance and reconstruction accuracy. Co-SLAM uses joint coordinate and sparse parametric scene for accurate scene representation. They can achieve promising reconstruction quality in a small indoor room.
@@
In local scene representation, We propose a parametric-coordinate joint encoding method for accuracy, speed, and completion of unseen region. Parametric encoding is the tri-plane encoding, and the coordinate encoding is the one-blob encoding with MLP. We use tri-planes to encode the local high-frequency feature of the scene and use MLP to represent global low-frequency features with the coherence priors inherent. We bring together the benefits of both methods for accuracy, smoothness, and hole-filling in areas without observation.
@@
  • A progressive scene representation method is proposed which dynamically initiate local scene representation trained with frames within a local window. This enables scalability to extensive indoor scenes and long videos sequences, substantially improving robustness.
  • In local scene representation, We design a joint parametric-coordinate encoding method. We combine the tri-planes with the one-blob encoding encoding method for accurate and smooth surface reconstruction. It can not only enhance the ability of scene representation, but also substantially reduce the memory growth from cubic to square.
  • We integrate the traditional SLAM system with an end-to-end pose estimation network. A local-to-global bundle adjustment algorithm is proposed, which can mitigate cumulative error in large-scale indoor scenes. Our system maintain a global keyframe database with the system operation, enabling bundle adjustment across all past observations, from local to global.
@@ 2 Related Work
Dense Visual SLAM. SLAM [15] and localization [16, 38, 37, 23] has been an active field for the past two decades. Traditional visual SLAM algorithms [19, 24, 9, 34, 35] estimate accurate camera poses and use sparse point clouds as the map representation. They use manipulated key points for tracking, mapping, relocalization, and loop closing. Dense visual SLAM approaches focus on reconstructing a dense map of a scene. DTAM [21] is one of the pioneer works that use the dense map and view-centric scene representation. KinectFusion [11] performs camera tracking via projective iterative-closest-point (ICP) and explicitly represents the surface of the environment via TSDF-Fusion. Some works [25, 5, 29] propose bundle adjustment(BA) method to optimize keyframe poses and construct the dense 3D structure jointly. In contrast to previous SLAM approaches, we adopt implicit scene representation of the geometry and directly optimize them during mapping.
Implicit Scene Representation.  With the proposal of Neural radiance fields (NeRF) [18], many researchers explore taking the advantages of the implicit method into 3D reconstruction. NeRF is a ground-breaking method for novel view synthesis using differentiable rendering. However, the representation of volume densities can not commit the geometric consistency. In order to deal with it, UNISURF [22] and NeuS [32] are proposed, combining neural radiance fields with Signed Distance Field (SDF) values. Other methods [1, 2, 3, 28] use various scene geometry representation methods, such as truncated signed distance function, voxel grid. For large-scale representation, Mega-NeRF and LocalRF [30, 17] use multiple local scene representations for the entire scene.
NeRF-based SLAM.  iMAP [27] and NICE-SLAM [41] are successively proposed to combine neural implicit mapping with SLAM. iMAP uses a single multi-layer perceptron (MLP) to represent the scene, and NICE-SLAM uses a learnable hierarchical feature grid. [40, 14, 13] use semantic feature embedding to improve scene representation. Some methods [7, 39] also use 3D gaussian to improve the scene representation. Co-SLAM [31] and ESLAM [12] are the most relative work of our method. However, all of them have difficult in large-scale indoor environments and long sequences. With the proposed progressive scene representation method, we can successfully scale up to larger indoor scenarios. The fusion of tri-planes and one-blob encoding leads to high-fidelity and smooth surface reconstruction in local scene. A local-to-global bundle adjustment method is also proposed. This method can effectively eliminate growing cumulative errors in existing methods in large indoor scenes.
@@ 3 Method
The pipeline of our system is shown in Fig. 1. We use a set of sequential RGB-D frames $\{I_{i},D_{i}\}_{i=1}^{M}$ with known camera intrinsic $K\in R_{3\times 3}$ as our input. Our model predicts camera poses $\{R_{i}|t_{i}\}^{M}_{i=1}$, color $\mathbf{c}$, and an implicit truncated signed distance function (TSDF) representation $\phi_{\boldsymbol{g}}$ that can be used in marching cubes algorithm to extract 3D meshes. For the implicit mapping thread, a progressive scene representation method (Sec. 3.1) is designed to represent large-scale indoor environments. Then, in the local radiance fields, we improve the scene representation methods and combine the tri-planes with multi-layer perceptron (MLP) by our designed architecture. Sec. 3.2 walks through the rendering process, which converts raw representations into pixel depths, colors, and SDF values. For the camera tracking thread, a local-to-global bundle adjustment method (Sec. 3.3) is designed for robust and accurate pose estimation. Several carefully designed loss functions are proposed to jointly optimize the scene implicit representation and camera pose estimation. The network is incrementally updated with the system operation.
@@ 3.1 Progressive Scene Representation
where $SR_{\theta_{n}}^{n}$ denotes the local scene representation, $\sigma$ denotes the volume density. Each local scene representation is centered at the position of the last estimated camera pose. We train each scene representation with a local subset frames. Each subset contains some overlap frames, which is important for achieving consistent reconstructions in the local scene representation. Whenever the estimated camera pose leaves the bound of the current scene representation, we stop optimizing previous ones (freeze the network parameters). At this point, we can reduce memory requirements by removing unnecessary supervisory frames. We also stop updating the mapping parameters in the tracking thread to reduce errors. If the estimated camera pose is outside the current bounds, but within a previous local scene representation, we activate the previous one and proceed with the optimization process. We further increase the global consistency by inverse distance weight (IDW) fusion for all overlapping scene representations at any supervising frame.
Local Scene Representation. Voxel grid-based architectures [10, 41, 31] are the mainstream in NeRF-based SLAM system. However, they struggle with cubical memory growing and real-time performance. Inspired by [12], we design a parametric-coordinate joint encoding method. Parametric encoding is tri-plane encoding, and the coordinate encoding is the one-blob encoding with MLP. We store and optimize high-frequency features(e.g. texture) on perpendicular axis-aligned planes. The one-blob encoding with MLPs are used to encode and store low-frequency features for the coherence and smoothness priors. This joint scene representation architecture achieve high-fidelity and smoothness scene reconstruction with the ability of hole filling.
Specifically, the tri-planes are at two scales, i.e., coarse and fine. The tri-planes feature $\boldsymbol{T}(x)$ can be formulated as:
@@
where $\gamma(\mathbf{x})$ represents coordinate position encoding. $\mathbf{z}$ is the latent code. We use one-blob encoding [31] instead of embedding spatial coordinates into multiple frequency bands. Finally, the color decoder predicts the RGB value:
@@
$\phi_{\boldsymbol{a}}(x)$ represents the color of the sample points. $\boldsymbol{a}(x)$ is the is the appearence feature from tri-planes. Combining the MLP with the tri-planes scene representation, our architecture achieve accurate and smooth surface reconstruction, efficient memory use, and hole filling performance.
@@ 3.2 Differentiable Rendering
where $\beta\in\mathbb{R}$ is a learnable parameter that controls the sharpness of the surface boundary. Then we define the termination probability $w_{k}$, depth $\hat{\boldsymbol{d}}$, and color $\hat{\boldsymbol{c}}$ as:
@@ 3.3 Local-to-global Bundle Adjustment
Currently, existing NeRF-based SLAM methods exhibit poor accuracy in large-scale indoor scene localization. Their tracking networks are performed via minimizing rgb loss functions with respect to learnable parameters $\theta$ to estimate the relative pose matrix $\{R_{i}|t_{i}\}\in\mathbb{SE}(3)$. With the growing cumulative error $\varepsilon$ of pose estimation, those methods result in failure in large-scale indoor scenes and long videos. To this end, we design a local-to-global bundle adjustment method to solve this problem, which performs well with our progressive scene representation. We design our method by drawing inspiration from traditional keyframe-based SLAM systems for improving the robustness and accuracy of pose estimation. We propose neural warpping error and reprojection error for local-to-global bundle adjustment. The neural warpping loss is formulated as:
@@
Whenever a keyframe arrives, we perform local bundle adjustment in our tracking and mapping thread. A keyframe is selected for every $K$ frames. When the camera moves to the bound of the current scene representation, we also initialize it as the keyframe. In local bundle adjustment, we only select keyframes from the local keyframe database that visually overlap with the current frame when optimizing the scene geometry to ensure the geometry outside the current view remains static and fast convergence. Meanwhile, we also maintain a global keyframe list with the operation of our system. After accumulating a specific number of local keyframes or the camera moves to the local bound, a global bundle adjustment is performed. In global BA, we randomly select keyframes and rays from the global keyframe database, which leverages all historical observations of the scene. This approach effectively integrates local and global information which greatly improves the accuracy of camera pose optimization in large-scale indoor scenes.
@@
Table 1: Quantitative results of our proposed PLGSLAM with existing NeRF-based SLAM system on the Replica dataset [26]. We evaluate reconstruction and localization performance in small room scenes. The results are the average on the scenes of the Replica dataset. Our method outperforms the existing method in surface reconstruction and pose estimation.
@@ 3.4 Objective Functions
Ours 6.033 11.086 3.261 85.357 5.574 6.228
Table 3: Quantitative results of our proposed PLGSLAM with existing NeRF-based SLAM system on the Apartment dataset [41]. We evaluate reconstruction and localization performance in large-scale multi-room scenes. The results are the average of three runs. Our method outperforms the existing method in surface reconstruction and pose estimation.
@@ 4.1 Datasets and Metrics
Figure 5: Completion ratio vs. model size and average time for PLGSLAM with other methods. Each model corresponds to a different hash-table or tri-planes size.
@@
Implementation We run PLGSLAM on a desktop PC with NVIDIA RTX 3090ti GPU. We employ feature planes with a resolution of 24 cm for coarse tri-planes. We use 6 cm resolution for fine tri-planes. All feature planes have 32 channels, resulting in a 64-channel concatenated feature input for the decoders. The decoders are two-layer MLPs with 32 channels in the hidden layer. For Replica [26], we sample $N$ = 32 points for stratified sampling and $N_{surface}$ = 8 points for importance sampling on each ray. And for ScanNet [4], we set $N$ = 48 and $N_{surface}$ = 8. For further details of our implementation, refer to the supplementary.